# ISP-AD: A Large-Scale Real-World Dataset for Advancing Industrial Anomaly Detection with Synthetic and Real Defects

### v2

@@ 6 Conclusions and Outlook
This publication introduces a novel surface defect detection dataset that exhibits a “real-world” industrial use case. The investigated screen-printed samples were captured using three different optical modalities. Typical for industrial inspection, the dataset reflects imperfect imaging conditions caused by imaging modalities, data preprocessing, and permitted sample variance, thereby presenting challenges for defect detection algorithms. In addition to extending common unsupervised settings, the dataset includes large-scale supervised training data, comprising tens of thousands of fault-free patches. Based on a review of recent literature, the proposed dataset can be assumed to be the largest publicly available, enabling research on both unsupervised and supervised approaches by using both synthetic and real defects.
It was observed that SOTA unsupervised methods, which achieve remarkable performance on established benchmark datasets such as MVTec [8], struggled in these specific industrial scenarios. The occurence of small and weakly contrasted defects, combined with high permitted design variability, significantly limited defect detection performance, especially at the pixel-levels. Among the evaluated methods, the most promising approaches are based on defect synthetization at both the feature and image-levels and knowledge distillation using patch descriptor networks, achieving image-level MCCs of up to 0.89 in LSM-1 and 0.96 in LSM-2. However, despite robust performance of Efficient-AD with industry applicable low FPRs in LSM-1 and LSM-2, weakly contrasted defects remained challenging. While certain unsupervised approaches have shown promising advances in image-level detection performance for specific imaging modalities (e.g., LSM-2), future research should focus enhancing segmentation performance under industrial imaging conditions.
Incorporating available weakly labeled defective samples within a mixed supervised training procedure is able to improve image-level defect detection performance by large margins. The experiments described in section 5.2 indicate that the initial comprehensive feature distribution, by means of several tens of thousands of synthesized punctual defects, is complemented and made more diverse, even with small amounts of available real defects. The injection of small fractions (9 samples in ASM) of previously unseen defect classes (area defects) substantially improved model generalization. The defect detection performance of small and weakly contrasted point defects could be significantly improved during mixed training (MCC $=$ 0.96 in LSM-1), compared to training on solely real defects (MCC $=$ 0.91).
@@
Furthermore, purely synthetic training achieved near-optimal (MCCs up to 0.99) defect detection performance in modalities that exhibit distinct feature characteristics, such as pinholes in LSM-2. Purely synthetic training outperformed SOTA unsupervised approaches in two out of three modalities (LSM-2 and ASM).
Despite promising investigations of generalization capabilities with a limited number of available defective samples, further experiments on more diverse defect classes with larger test sets are necessary to gain deeper insights.
@@
Additionally, permitted normal data variations are learned through large amounts (50 0005000050\,00050 000100 000100000100\,000100 000 of augmented fault-free patches, even with noisy labels. Noisy labels, due to e.g. process contamination during the image acquisition, were estimated to account for 2 – 5 % in LSM-1. These corrupted labels did not show a significant impact during training. Thus, the screening effort of fault-free data in the proposed unsupervised training datasets could be considered more elaborate than the presented supervised data preprocessing workflow in [35], minimizing manual labeling effort to known defective sample regions. Thus, further research should focus on efficient and industrial applicable data preparation and preprocessing workflows within unsupervised and supervised approaches and the influence of different proportions of noisy labels.
As a direct implication of above findings, research in self-supervised approaches that utilize these representative normal data distributions should be emphasized to enhance industry-applicable model robustness at low labeling costs.
Recent supervised anomaly detection methods [57, 76, 34] align with our findings, demonstrating improved detection performance through the integration of limited labeled data in the mixed training process. Those results indicate that incorporating both synthetic and real defects during training is a promising strategy in industrial defect detection.
@@ 6.1 Future Applications
  • Investigation of the impact of increased normal feature variance on the robustness of unsupervised approaches by utilizing larger sets of fault-free training samples than those used in the proposed experiments. Additionally, an examination of the effect of real noisy labels by applying the “unscreened” automatically extracted fault-free training data.
  • Conduction of research on self-supervised learning approaches [29] utilizing the synthetic or fault-free large scale training sets. Therefore, feature extractors of unsupervised approaches, such as described in [33, 41], could be pretrained to learn more discriminative features of the target domain.
@@
Data availability The dataset supporting this research is publicly available at Zenodo https://doi.org/10.5281/zenodo.14911043.
@@
Modality Good Good #Synthetic #Area #Points Good #Area #Points
LSM-1 3678 100 669100669100\,669100 669 99 3319933199\,33199 331 120 52 1470 49 46
LSM-2 401 50 0365003650\,03650 036 49 9644996449\,96449 964 62 66 382 63 33
ASM 500 103 622103622103\,622103 622 96 3699636996\,36996 369 37 109 1916 33 41
@@
Table 7 illustrates the training and test data splits of the proposed ISP-AD dataset. The augmentation settings applied to the supervised training data are shown in Table 8. In addition to the presented data splits, large-scale fault-free training data (10 $0 0$ 20 0002000020\,00020 000 patches) without augmentation is made available for all three modalities: LSM-1, LSM-2, and ASM. As outlined in Section 3.2, the extracted patches can be considered noisy. The naming convention of image-level labels within the .hdf5 files follows:

### v3

@@ 6 Conclusions and Outlook
This publication introduces a novel surface defect detection dataset that exhibits a “real-world” industrial use case. The investigated screen-printed samples were captured using three different optical modalities. Typical for industrial inspection, the dataset reflects imperfect imaging conditions caused by imaging modalities, data preprocessing, and permitted sample variations, thereby presenting challenges for defect detection algorithms. In addition to extending common unsupervised settings, the dataset includes large-scale supervised training data, comprising tens of thousands of fault-free patches. Based on a review of recent literature, the proposed dataset can be assumed to be the largest publicly available, enabling research on both unsupervised and supervised approaches by using both synthetic and real defects.
It was observed that SOTA unsupervised methods, which achieve remarkable performance on established benchmark datasets such as MVTec [9], struggled in these specific industrial scenarios. The occurrence of small and weakly contrasted defects, combined with high permitted design variability, significantly limited defect detection performance, especially at the pixel level. Among the evaluated methods, the most promising approaches are based on defect synthetization at both the feature and image-levels and knowledge distillation using patch descriptor networks, achieving image-level MCCs of up to 0.89 in LSM-1 and 0.96 in LSM-2. However, despite robust performance of Efficient-AD [7] with industry-applicable low FPRs in LSM-1 and LSM-2, weakly contrasted defects remained challenging. While certain unsupervised approaches have shown promising advances in image-level detection performance for specific imaging modalities (e.g., LSM-2), future research should focus enhancing segmentation performance under industrial imaging conditions.
Incorporating available weakly labeled defective samples within a mixed supervised training strategy is able to improve image-level defect detection performance by large margins. The experiments described in Section 5.2 indicate that the initial comprehensive feature distribution, by means of several tens of thousands of synthesized punctual defects, is complemented and made more diverse, even with small amounts of available real defects. The injection of small fractions (9 samples in ASM) of previously unseen defect classes (area defects) substantially improved model generalization. These results indicate that previously unseen defects, such as those that could not be synthesized, can provide strong learning signals, helping to refine the decision boundary.
The defect detection performance of small and weakly contrasted point defects was substantially improved with mixed training (MCC = 0.96 in LSM-1), compared to training solely on real defects (MCC = 0.91) using a standard ResNet-18 backbone. Notably, in the few-shot regime, the mixed supervised strategy showed substantial performance gains compared to pure oversampling across various backbones (ResNet-18, ResNet-50, EfficientNet-B0, EfficientNet-B4, and ConvNeXt-Tiny). More recent backbones, such as ConvNeXt-Tiny, achieved a high MCC of up to 0.96 with only 10 real injected defects.
@@
Furthermore, purely synthetic training achieved near-optimal (MCCs up to 0.99) defect detection performance in modalities that exhibit distinct feature characteristics, such as pinholes in LSM-2. Purely synthetic training outperformed SOTA unsupervised approaches in two out of three modalities (LSM-2 and ASM; ResNet-18).
Despite promising investigations of generalization capabilities with a limited number of available defective samples, further experiments are necessary for deeper insights. These should involve more diverse defect classes and varying hyperparameters, such as the injection probability $p_{\text{inj}}$ of real defects. The selected injection probability of 1/32 was intended to avoid overfitting to injected real defect characteristics. However, in some cases—depending on defect availability and diversity, the defect synthesis approach, or the imaging modality (e.g., ASM)—higher injection probabilities may be preferable to align the decision boundary more closely with the incorporated defect features. Furthermore, in future work, defect selection strategies inspired by active learning frameworks [25] could be considered, as higher fractions of real defects appear to saturate performance in certain cases. Therefore, injected diversity seems to be more important than the absolute quantity of defective data.
Additionally, permitted normal data variations are learned through large amounts (50 00050\,000100 000100\,000 of augmented fault-free patches, even with noisy labels. Noisy labels, due to e.g. process contamination during the image acquisition, were estimated to account for 2 – 5 % in LSM-1. These corrupted labels did not show a significant impact during training. This observation is consistent with unsupervised and supervised learning paradigms: unsupervised methods rely on clean normal data to enable the modeling of compact representations, whereas supervised methods are able to learn a robust decision boundary even in the presence of noisy labels [6]. Regarding our approach, this can be explained by the large-scale balanced data streams, including a diverse set of permitted fault-free data variations, helping the model to learn what patterns are considered normal.
@@
Thus, the screening effort of fault-free data in the proposed unsupervised training datasets could be considered more elaborate than the presented supervised data preprocessing workflow in [38], minimizing manual labeling effort to known defective sample regions. Thus, further research should focus on efficient and industrial applicable data preparation and preprocessing workflows within unsupervised and supervised approaches and the influence of different proportions of noisy labels. As a direct implication of above findings, research in self-supervised approaches that utilize these large-scale normal data distributions should be emphasized to enhance industry-applicable model robustness at low labeling costs.
In summary, the mixed supervised training strategy [28], based on the efficient preprocessing workflow proposed in [38], is capable of fulfilling industrial inspection requirements with minimized labeling effort. Labeling is restricted to known defective samples and regions, as the extraction of fault-free patches is integrated into an automated procedure. The model-free synthetization algorithm enables defect generation without requiring elaborate learning of defect distributions. Complementary defect features are learned through the injection of a small set of weakly labeled real defects, enabling model generalization. However, recent zero-shot and few-shot anomaly synthesis approaches [69, 70, 41] may be preferable in future adaptations, further enhancing defect realism and diversity within the synthetic baseline.
Based on our findings, we raise the question of whether defining industrial anomaly detection as a fully open-set problem is appropriate in common real-world manufacturing processes. Since defect causes are often known in advance, a reformulated paradigm such as the proposed restricted open-set problem may offer a more practically applicable definition for industrial scenarios. Future work may explore this direction in more detail, particularly in terms of formalizing such paradigms and resulting learning strategies, and evaluating them under realistic industrial constraints.
@@ 6.1 Future Applications
  • Investigation of the impact of increased normal feature variations on the robustness of unsupervised approaches by utilizing larger sets of fault-free training samples than those used in the proposed experiments. Additionally, an examination of the effect of real noisy labels by applying the “unscreened” automatically extracted fault-free training data.
  • Conduction of research on self-supervised learning approaches [31] utilizing the synthetic or fault-free large-scale training sets. Therefore, feature extractors of unsupervised approaches, such as described in [36, 45], could be pretrained to learn more discriminative features of the target domain.
  • Foster research on zero-shot or few-shot defect synthesis approaches like recent works [88, 69, 70, 41] that may enable rapid initial deployment in industrial scenarios. Therefore, the additional fault-free data in ISP-AD can serve as a research basis.
@@
Data availability The dataset supporting this research is publicly available on Zenodo at https://doi.org/10.5281/zenodo.14911042, and is licensed under the Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International (CC BY-NC-SA 4.0).
@@
Modality #Good #Good #Synthetic #Area #Points #Good #Area #Points
LSM-1 3678 100 669100\,669 99 33199\,331 120 52 1470 49 46
LSM-2 401 50 03650\,036 49 96449\,964 62 66 382 63 33
ASM 500 103 622103\,622 96 36996\,369 37 109 1916 33 41
@@
Table 7 illustrates the training and test data splits of the proposed ISP-AD dataset. The augmentation settings applied to the supervised training data are shown in Table 8. In addition to the presented data splits, large-scale fault-free training data (10 00010\,000 - 20 00020\,000 patches) without augmentation is made available for all three modalities: LSM-1, LSM-2, and ASM. As outlined in Section 3.2, the extracted patches can be considered noisy. The naming convention of image-level labels within the .hdf5 files follows: